# Absence of antisymmetric tensor fields: Clue from Starobinsky model of $f(R)$ gravity

### v3

@@ 1 Introduction
It was shown earlier that the presence of warped extra spatial dimension may lead to suppression of antisymmetric tensor field of various ranks[7, 8, 9, 10, 11]. While any extra-dimensional theory brings in a plethora of other predictions hitherto unobserved, within the domain of 3+ 1 dimensional Einstein gravity , it is hard to explain the reason of the disappearance of the massless KR field in our observable universe. As an alternative, in this work, we want to explore a dynamical mechanism of this suppression of the KR field as well as the other higher rank anti-symmetric fields in the light of higher curvature $f(R)$ theory. The presence of suitable higher curvature terms in the gravity action is permissible as long as they satisfy diffeomorphic invariance [12, 13, 14, 15, 16]. In this context $f(R)$ theories invoked great interest in the context of various cosmological predictions so far unexplained[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Higher order curvature terms in the gravity actions in such models are automatically suppressed by Planck scale which explains why at low energy only Einstein’s term dominates over the others. However, at high energy or at the epoch of early universe , the higher curvature term will have significant presence and in this context Starobinsky model [29] has been studied extensively in different context[30, 31, 32, 33, 34, 35, 36]. In this work, we explore whether the dynamical evolution in the presence of higher curvature gravity in 3+1 dimension may lead to a natural geometric suppression of the massless higher rank antisymmetric tensor fields.
In section 2, we present the standard formalism of transformation of an action in $f(R)$ gravity to Einstein gravity through a conformal transformation which manifests the higher curvature in terms of a scalar degree of freedom along with a potential. This transformation has been extensively discussed in the literature[12, 13, 17]. In section 3, we present the field equations of the transformed action in an isotropic and homogeneous universe. The conformal transformation induces coupling of scalar field with different matter. The coupled nonlinear equations are presented as a set of autonomous system in terms of a new set of dynamical variables. The form of $f(R)$ and the potentials generated from this modified gravity are discussed in section 4. Section 5 presents the solutions of the field equations in three different cases - A) $n=2$ or the Starobinsky model, B) $n\neq 2$ or generalised Starobinsky model and C) Both the models with Cosmological constant. In section 6 we highlight the finding of our work and finally, we discuss and conclude in section 7.
@@ 2 $f(R)$ Formalism and Conformal Transformation
For a rank $n$ antisymmetric tensor field $X_{a_{1}a_{2}...a_{n}}$, the field strength tensor of the massless mode can be written as $Y_{a_{1}a_{2}...a_{n+1}}=\partial_{[a_{n+1}}X_{a_{1}a_{2}...a_{n}]}$. In 3+1 dimensional spacetime, the maximum rank of an antisymmtric tensor field can be 3 as the field strength vanishes beyond that. The action of higher curvature gravity in 3+1 dimension along with other fields can be written as
$\begin{split}S_{\text{gravity}}=\int d^{4}x\sqrt{-g}\bigg{(}\frac{F(R)}{2\kappa^{2}}+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\bigg{)}\end{split}$ (2.1)
@@
It is standard to represent [12, 13] the action (2.1) in the following way with an introduction of an auxiliary field
$S=$ $\int d^{4}x\sqrt{-g}\Bigg{(}\frac{1}{2\kappa^{2}}\phi R-V(\phi)+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\Bigg{)}$ (2.2)
@@
$\phi\equiv f^{\prime}(R)~{}\text{and}~{}V(\phi)=\frac{1}{2\kappa^{2}}[{\phi~{}R-f(R)}]$ (2.3)
Now, following a conformal transformation of the form $\tilde{g}_{\mu\nu}(x)=\Omega^{2}{g}_{\mu\nu}(x)$, where $\Omega^{2}=f^{\prime}=\phi$ one can rewrite the action (2.2) in the Einstein frame as
$S_{E}=$ $\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})+e^{\sqrt{\frac{2}{3}}k\tilde{\phi}}\Psi^{+}\tilde{\gamma}^{0}i\tilde{\gamma}^{\mu}D_{\mu}{\Psi}-\frac{1}{4}\widetilde{g}^{\mu\alpha}\widetilde{g}^{\nu\beta}{F}_{\mu\nu}{F}_{\alpha\beta}$ (2.4)
@@ 3 Field Equations in Einstein Frame
$S_{E}$ $=\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})\bigg{)}$ (3.1)
@@
${\tilde{T}^{\tilde{\phi}}}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial(\sqrt{-\tilde{g}}{{\cal L}_{\tilde{\phi}}})}{\partial{\tilde{g}}^{\mu\nu}}~{}~{}\text{and}~{}~{}{\tilde{T}}^{M}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{g}}^{\mu\nu}}$ (3.2)
Variation of the action (3.1) with respect to the field $\tilde{\phi}$ gives the dynamical equation for the scalar field
$\square{\tilde{\phi}}-\frac{\partial V}{\partial{\tilde{\phi}}}+\frac{1}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=0$ (3.3)
@@
$\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=\frac{\partial{\cal L}_{M}}{\partial g^{\mu\nu}}\frac{\partial g^{\mu\nu}}{\partial{\tilde{\phi}}}={\sqrt{-\tilde{g}}}\kappa Q{\tilde{T}}$ (3.4)
@@
The matter in (3.2) is considered in the form of perfect fluid $\tilde{T}^{\mu(M)}_{\nu}=\text{diag}(-\tilde{\rho}_{M},\tilde{P}_{M},\tilde{P}_{M},\tilde{P}_{M})$. $\tilde{\rho}_{M}$ and $\tilde{P}_{M}$ are the energy density and pressure of matter. We consider the observable universe to consist of radiation and matter (baryonic and CDM) i.e., $\tilde{\rho}_{M}=\tilde{\rho}_{r}+\tilde{\rho}_{m}$.
@@
$3\tilde{H}^{2}=\kappa^{2}[\frac{1}{2}(\frac{d\tilde{\phi}}{d\tilde{t}})^{2}+V(\tilde{\phi})+\tilde{\rho}_{M}]$ (3.5)
$\frac{d^{2}\tilde{\phi}}{d\tilde{t}^{2}}+3\tilde{H}\frac{d\tilde{\phi}}{d\tilde{t}}+\frac{\partial V(\tilde{\phi})}{\partial{\tilde{\phi}}}=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})$ (3.6)
$\frac{d\tilde{\rho}_{M}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{M}+\tilde{P}_{M})=\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (3.7)
where $\tilde{H}\equiv\frac{1}{\tilde{a}}\frac{d\tilde{a}}{\tilde{d}t}$. One can rewrite equation (3.6) as
$\frac{d\tilde{\rho}_{\phi}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (3.8)
where $\tilde{\rho}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})+V(\tilde{\phi}))$ is the energy density of the scalar field and $\tilde{P}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})-V(\tilde{\phi}))$ is the pressure of the scalar field Using the standard thermodynamic relation $\tilde{P}_{M}=\omega\tilde{\rho}_{M}$, one can easily find a relation between $\tilde{\rho}_{M}$ and $\tilde{\phi}$ from equation (3.7)
$\tilde{\rho}_{M}=\tilde{\rho}_{M0}\tilde{a}^{-3(1+\omega)}\exp({\frac{\kappa}{\sqrt{6}}(1-3\omega)(\tilde{\phi}-\tilde{\phi}_{0})})$ (3.9)
@@
$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}\bigg{]}$ (3.10)
$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m})\tilde{\phi}^{\prime}$ (3.11)
$\tilde{\rho}_{m}^{\prime}+3\tilde{\rho}_{m}=\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{m}\tilde{\phi}^{\prime}$ (3.12)
$\tilde{\rho}_{r}^{\prime}+4\tilde{\rho}_{r}=0$ (3.13)
where prime denotes the differentiation with respect to $N(=\ln{\tilde{a}})$. Note that since the radiation field is traceless the scalar field does not couple to radiation (3.4). In order to solve these coupled nonlinear inhomogeneous equations, we introduce a set of dynamical variables [37, 38, 39]
$x=\frac{\kappa\tilde{\phi}\prime}{\sqrt{6}},y=\frac{\kappa\sqrt{V}}{\sqrt{3}\tilde{H}},\lambda=-\frac{V_{,\tilde{\phi}}}{kV},~{}m=\frac{\kappa^{2}\tilde{\rho}_{m}}{3\tilde{H}^{2}},~{}r=\frac{\kappa^{2}\tilde{\rho}_{r}}{3\tilde{H}^{2}}$ (3.14)
and rewrite the equations (3.10-3.13) in terms of these new variables forming a set of autonomous systems of first-order differential equations
@@
where, $\Gamma=\frac{VV_{\tilde{\phi}\tilde{\phi}}}{V_{\tilde{\phi}}^{2}}$. In terms of these new variables, equation (3.10) takes the form $1=x^{2}+y^{2}+m+r$.
@@ 4 General class of $f(R)$ : $f(R)=R+\alpha_{n}R^{n}$
In this work, we consider a very general class of modified gravity $f(R)=R+\alpha_{n}R^{n}$, where $\alpha_{n}$ has a suitable dimension. The Starobinsky model of modified gravity corresponds to $n=2$ i.e., $f(R)=R+\alpha_{2}R^{2}$, where $\alpha_{2}\sim\frac{1}{M^{2}}$. With this general form of $f(R)$, the corresponding potential is
$V(\tilde{\phi})={\cal A}e^{-2\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}\bigg{(}e^{\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}-1\bigg{)}^{\frac{n}{n-1}}$ (4.1)
where ${\cal A}=\frac{(n-1)\alpha_{n}}{2\kappa^{2}}\big{[}\frac{1}{n\alpha_{n}}\big{]}^{\frac{n}{n-1}}$. The potential becomes zero at $\tilde{\phi}=0$, irrespective of $n$. For $n=2$, the potential is constant when $\tilde{\phi}$ is positive. For negative $\tilde{\phi}$ it increases exponentially. For $n\neq 2$, the potential increases sharply for negative $\tilde{\phi}$ and goes to 00 for positive $\tilde{\phi}$. In fig (1) we present the form of potentials for various values of $n$ with ${\cal A}=1$.

### v4

@@ Abstract
keywords:
f(R) Theory, Einstein Frame, Dynamical Dark Energy, Torsion
journal: Physics of the Dark Universe
[…]
@@ 1 Introduction
It was shown earlier that the presence of warped extra spatial dimension may lead to suppression of antisymmetric tensor field of various ranks[7, 8, 9, 10, 11]. While any extra-dimensional theory brings in a plethora of other predictions hitherto unobserved, within the domain of 3+ 1 dimensional Einstein gravity , it is hard to explain the reason of the disappearance of the massless KR field in our observable universe. As an alternative, in this work, we want to explore a dynamical mechanism of this suppression of the KR field as well as the other higher rank anti-symmetric fields in the light of higher curvature $f(R)$ theory. The presence of suitable higher curvature terms in the gravity action is permissible as long as they satisfy diffeomorphic invariance [12, 13, 14, 15, 16]. In this context $f(R)$ theories invoked great interest in the context of various cosmological predictions so far unexplained[17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. Higher order curvature terms in the gravity actions in such models are automatically suppressed by Planck scale which explains why at low energy only Einstein’s term dominates over the others. However, at high energy or at the epoch of early universe , the higher curvature term will have significant presence and in this context Starobinsky model [29] has been studied extensively in different context[30, 31, 32, 33, 34, 35, 36]. In this work, we explore whether the dynamical evolution of the components in the presence of higher curvature gravity in 3+1 dimension may lead to a natural geometric suppression of the massless higher rank antisymmetric tensor fields.
In section 2, we present the standard formalism of transformation of an action in $f(R)$ gravity to Einstein gravity through a conformal transformation which manifests the higher curvature in terms of a scalar degree of freedom along with a potential. This transformation has been extensively discussed in the literature[12, 13, 17]. In section 3, we present the field equations of the transformed action in an isotropic and homogeneous universe. The conformal transformation induces coupling of scalar field with different component. The coupled nonlinear equations are presented as a set of autonomous system in terms of a new set of dynamical variables. The form of $f(R)$ and the potentials generated from this modified gravity are discussed in section 4. Section 5 presents the solutions of the field equations in three different cases - A) $n=2$ or the Starobinsky model, B) $n\neq 2$ or generalised Starobinsky model and C) Both the models with Cosmological constant. In section 6 we highlight the finding of our work and finally, we discuss and conclude in section 7.
@@ 2 $f(R)$ Formalism and Conformal Transformation
For a rank $n$ antisymmetric tensor field $X_{a_{1}a_{2}...a_{n}}$, the field strength tensor of the massless mode can be written as $Y_{a_{1}a_{2}...a_{n+1}}=\partial_{[a_{n+1}}X_{a_{1}a_{2}...a_{n}]}$. In 3+1 dimensional spacetime, the maximum rank of an antisymmtric tensor field can be 3 as the field strength vanishes beyond that. In an early universe post reheating epoch (around the time of nucleosynthesis), the effective action of higher curvature gravity in 3+1 dimension along with other fields can be written as
$\begin{split}S_{\text{gravity}}=\int d^{4}x\sqrt{-g}\bigg{(}\frac{F(R)}{2\kappa^{2}}+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\bigg{)}\end{split}$ (1)
@@
It is standard to represent [12, 13] the action (1) in the following way with an introduction of an auxiliary field
$S=$ $\int d^{4}x\sqrt{-g}\Bigg{(}\frac{1}{2\kappa^{2}}\phi R-V(\phi)+\bar{\Psi}i\gamma^{\mu}D_{\mu}\Psi-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{4}H_{\mu\nu\rho}H^{\mu\nu\rho}-\frac{1}{4}T_{\mu\nu\rho\delta}T^{\mu\nu\rho\delta}\Bigg{)}$ (2)
@@
$\phi\equiv f^{\prime}(R)~{}\text{and}~{}V(\phi)=\frac{1}{2\kappa^{2}}[{\phi~{}R-f(R)}]$ (3)
Now, following a conformal transformation of the form $\tilde{g}_{\mu\nu}(x)=\Omega^{2}{g}_{\mu\nu}(x)$, where $\Omega^{2}=f^{\prime}=\phi$ one can rewrite the action (2) in the Einstein frame as
$S_{E}=$ $\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})+e^{\sqrt{\frac{2}{3}}k\tilde{\phi}}\Psi^{+}\tilde{\gamma}^{0}i\tilde{\gamma}^{\mu}D_{\mu}{\Psi}-\frac{1}{4}\widetilde{g}^{\mu\alpha}\widetilde{g}^{\nu\beta}{F}_{\mu\nu}{F}_{\alpha\beta}$ (4)
@@ 3 Field Equations in Einstein Frame
$S_{E}$ $=\int d^{4}x\sqrt{-\tilde{g}}\bigg{(}\frac{\tilde{R}}{2\kappa^{2}}-\frac{1}{2}\tilde{g}^{\mu\nu}\partial_{\mu}{\tilde{\phi}}\partial_{\nu}{\tilde{\phi}}-V(\tilde{\phi})\bigg{)}$ (5)
@@
${\tilde{T}^{\tilde{\phi}}}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial(\sqrt{-\tilde{g}}{{\cal L}_{\tilde{\phi}}})}{\partial{\tilde{g}}^{\mu\nu}}~{}~{}\text{and}~{}~{}{\tilde{T}}^{M}_{\mu\nu}=-\frac{2}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{g}}^{\mu\nu}}$ (6)
Variation of the action (5) with respect to the field $\tilde{\phi}$ gives the dynamical equation for the scalar field
$\square{\tilde{\phi}}-\frac{\partial V}{\partial{\tilde{\phi}}}+\frac{1}{\sqrt{-\tilde{g}}}\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=0$ (7)
@@
$\frac{\partial{\cal L}_{M}}{\partial{\tilde{\phi}}}=\frac{\partial{\cal L}_{M}}{\partial g^{\mu\nu}}\frac{\partial g^{\mu\nu}}{\partial{\tilde{\phi}}}={\sqrt{-\tilde{g}}}\kappa Q{\tilde{T}}$ (8)
@@
The matter in (6) is considered in the form of perfect fluid $\tilde{T}^{\mu(M)}_{\nu}=\text{diag}(-\tilde{\rho}_{M},\tilde{P}_{M},\tilde{P}_{M},\tilde{P}_{M})$. $\tilde{\rho}_{M}$ and $\tilde{P}_{M}$ are the energy density and pressure of matter. We consider the observable universe to consist of radiation and matter (baryonic and CDM) i.e., $\tilde{\rho}_{M}=\tilde{\rho}_{r}+\tilde{\rho}_{m}$.
@@
$3\tilde{H}^{2}=\kappa^{2}[\frac{1}{2}(\frac{d\tilde{\phi}}{d\tilde{t}})^{2}+V(\tilde{\phi})+\tilde{\rho}_{M}]$ (9)
$\frac{d^{2}\tilde{\phi}}{d\tilde{t}^{2}}+3\tilde{H}\frac{d\tilde{\phi}}{d\tilde{t}}+\frac{\partial V(\tilde{\phi})}{\partial{\tilde{\phi}}}=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})$ (10)
$\frac{d\tilde{\rho}_{M}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{M}+\tilde{P}_{M})=\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (11)
where $\tilde{H}\equiv\frac{1}{\tilde{a}}\frac{d\tilde{a}}{\tilde{d}t}$. One can rewrite equation (10) as
$\frac{d\tilde{\rho}_{\phi}}{d\tilde{t}}+3\tilde{H}(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\kappa Q(\tilde{\rho}_{M}-3\tilde{P}_{M})\frac{d\tilde{\phi}}{d\tilde{t}}$ (12)
where $\tilde{\rho}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})+V(\tilde{\phi}))$ is the energy density of the scalar field and $\tilde{P}_{\tilde{\phi}}=(\frac{1}{2}(\frac{d^{2}\tilde{\phi}}{\tilde{d}t^{2}})-V(\tilde{\phi}))$ is the pressure of the scalar field Using the standard thermodynamic relation $\tilde{P}_{M}=\omega\tilde{\rho}_{M}$, one can easily find a relation between $\tilde{\rho}_{M}$ and $\tilde{\phi}$ from equation (11)
$\tilde{\rho}_{M}=\tilde{\rho}_{M0}\tilde{a}^{-3(1+\omega)}\exp({\frac{\kappa}{\sqrt{6}}(1-3\omega)(\tilde{\phi}-\tilde{\phi}_{0})})$ (13)
@@
$3\tilde{H}^{2}=\kappa^{2}\bigg{[}\frac{1}{2}{\tilde{\phi}^{\prime 2}}\tilde{H}^{2}+V(\tilde{\phi})+\tilde{\rho}_{m}+\tilde{\rho}_{r}\bigg{]}$ (14)
$\tilde{\rho}_{\phi}^{\prime}+3(\tilde{\rho}_{\tilde{\phi}}+\tilde{P}_{\tilde{\phi}})=-\frac{\kappa}{\sqrt{6}}(\tilde{\rho}_{m})\tilde{\phi}^{\prime}$ (15)
$\tilde{\rho}_{m}^{\prime}+3\tilde{\rho}_{m}=\frac{\kappa}{\sqrt{6}}\tilde{\rho}_{m}\tilde{\phi}^{\prime}$ (16)
$\tilde{\rho}_{r}^{\prime}+4\tilde{\rho}_{r}=0$ (17)
where prime denotes the differentiation with respect to $N(=\ln{\tilde{a}})$. Note that since the radiation field is traceless the scalar field does not couple to radiation (8). In order to solve these coupled nonlinear inhomogeneous equations, we introduce a set of dynamical variables [37, 38, 39]
$x=\frac{\kappa\tilde{\phi}\prime}{\sqrt{6}},y=\frac{\kappa\sqrt{V}}{\sqrt{3}\tilde{H}},\lambda=-\frac{V_{,\tilde{\phi}}}{kV},~{}m=\frac{\kappa^{2}\tilde{\rho}_{m}}{3\tilde{H}^{2}},~{}r=\frac{\kappa^{2}\tilde{\rho}_{r}}{3\tilde{H}^{2}}$ (18)
and rewrite the equations (14-17) in terms of these new variables forming a set of autonomous systems of first-order differential equations
@@
where, $\Gamma=\frac{VV_{\tilde{\phi}\tilde{\phi}}}{V_{\tilde{\phi}}^{2}}$. In terms of these new variables, equation (14) takes the form $1=x^{2}+y^{2}+m+r$.
@@ 4 General class of $f(R)$ : $f(R)=R+\alpha_{n}R^{n}$
In this work, we consider a very general class of modified gravity $f(R)=R+\alpha_{n}R^{n}$, where $\alpha_{n}$ has a suitable dimension. The Starobinsky model of modified gravity corresponds to $n=2$ i.e., $f(R)=R+\alpha_{2}R^{2}$, where $\alpha_{2}\sim\frac{1}{M^{2}}$, where $M$ is in MeV scale. With this general form of $f(R)$, the corresponding potential is
$V(\tilde{\phi})={\cal A}e^{-2\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}\bigg{(}e^{\sqrt{\frac{2}{3}}\kappa\tilde{\phi}}-1\bigg{)}^{\frac{n}{n-1}}$ (20)
where ${\cal A}=\frac{(n-1)\alpha_{n}}{2\kappa^{2}}\big{[}\frac{1}{n\alpha_{n}}\big{]}^{\frac{n}{n-1}}$. The potential becomes zero at $\tilde{\phi}=0$, irrespective of $n$. For $n=2$, the potential is constant when $\tilde{\phi}$ is positive. For negative $\tilde{\phi}$ it increases exponentially. For $n\neq 2$, the potential increases sharply for negative $\tilde{\phi}$ and goes to 00 for positive $\tilde{\phi}$. In fig (1) we present the form of potentials for various values of $n$. The form of the potentials are independent of the choice of $\cal A$.